A train starts from a station nd travels with uniform acceleration 0.5ms for 20 second. it travels with uniform velocity for another 30 second , the brakes are then applied so that a uniform retardation is obtained and the train comes to rest in a further 10 second. Sketch the velocity time graph of this motion. Using ur graph, Calculate the total distance travelled by the train.


1
Expert's answer
2021-11-01T19:25:01-0400



"v=a_1t_1=0.5\cdot 20=10~\frac ms," "s=s_1+s_2+s_3=\frac{vt_1}2+vt_2+\frac{vt_3}2=100+300+50=450~m."